[2+3) power 2 - 3 .5 +7
thank you sarah

Answer by fractalier(2101)

Okay from
(2 + 3)^2 - 3 * 5 + 7
we apply our order of operations rules...PEMDAS...
(2 + 3)^2 - 3 * 5 + 7 =
(5)^2 - 3 * 5 + 7 =
25 - 3 * 5 + 7 =
25 - 15 + 7 =
10 + 7 =
17